OpenStudy (anonymous):

10.5 • (–3) ÷ 10 • (–2) ÷ 0.5 simplify

OpenStudy (anonymous):

Use Operations order to simplify .

OpenStudy (nathan917):

10.5/(-3)=-31.5 -31.5/10=-3.15 -3.15*-2= 6.3 6.3/0.5=12.6

OpenStudy (anonymous):

thank you nathan

OpenStudy (nathan917):

Anytime.